how do you solve the equation by using the square roots: 3x^2-27=0
Answer by Alan3354(69443)

how do you solve the equation by using the square roots: 3x^2-27=0
-------------
Divide by 3
x^2-9 = 0
x^2=9
x = 3
x = -3